# Gaussian-Process-based Adaptive Tracking Control with Dynamic Active Learning for Autonomous Ground Vehicles

### v1

@@ II-C Online Learning and Update for SGPs
The VFE method is constructed for offline training, i.e., it requires the complete training set to be available. For adaptive control, online training is more suitable as it updates the model during operation using the latest measurement data. For training GPs online, various algorithms such as the recursive least squares GP (RLS-GP) [11] or the dynamic sparse GP (DS-GP) [10] are available. This article proposes a recursive gradient-based (RGB) optimization scheme for SGPs utilizing the VFE method outlined in Sec.  II-B. Compared to the previously introduced techniques that either update the approximate posterior distribution (RLS-GP) or the inducing points (DS-GP), this approach jointly updates both the hyperparameters and the inducing points of the GP based on the incoming new information, which to the authors knowledge has not been applied before.
First, we assume that the SGP has been trained offline on the initial dataset $\mathcal{D}^{N}$, yielding the optimal values for $\theta$ hyperparameters and $X^{M}$ inducing points that correspond to $Y^{M}$ virtual outputs. As outlined before, the core idea of sparse GP estimation is that the virtual $D^{M}$ dataset is obtained such that its distribution approximates the original $D^{N}$ set with considerably fewer data points. Therefore, in future predictions, we can rely only on the $D^{M}$ set, as it represents the information content of $\mathcal{D}^{N}$.
Let the one incoming batch of input-output data at time step $k$ be denoted as $\mathcal{D}^{Z}_{k}=\{X^{Z}_{k},Y^{Z}_{k}\}$, where $X^{Z}_{k}=[x_{1}^{\top}\cdots x_{Z}^{\top}]\in\mathbb{R}^{Z\times n_{x}}$, $Y^{Z}_{k}=[y_{1}\cdots y_{Z}]\in\mathbb{R}^{Z}$ and $Z$ is the batch size. Then, after receiving one batch of new training data at time $k$, we can combine the previously obtained pseudo dataset and the new batch to define a new training dataset as $\mathcal{\hat{D}}_{k}=\mathcal{D}^{M}_{k-1}\cup\mathcal{D}^{Z}_{k}$, where $\mathcal{D}_{0}^{M}=\mathcal{D}^{M}$. Note that with this formulation, $\mathcal{\hat{D}}_{k}$ both contains information from the pre-trained GP and innovation from the update batch. By substituting this combined dataset into the VFE cost (8), we can perform $n_{\alpha}$ steps of gradient descent to collectively update the hyperparameters of the GP and the inducing points, i.e.
@@
where $\alpha$ is the learning rate and $i\in\mathbb{I}_{1}^{n_{\alpha}}$. Note that the evaluation of (11) has the computational complexity of $\mathcal{O}((M+Z)M^{2})$, therefore, gradient descent can be executed multiple times online in one update iteration. Furthermore, using the augmented training dataset $\mathcal{\hat{D}}_{k}$ is beneficial as it incorporates both the previously learned knowledge from $D^{M}_{k-1}$ and the new information contained in $\mathcal{D}_{k}^{Z}$. The RBG algorithm is outlined in Alg. 1 and has three important parameters: the size of the pseudo dataset $M$ and the update batch size $Z$ and $n_{\alpha}$ number of retrain iterations.
@@
2:  $\mathcal{\hat{D}}_{k}\leftarrow\mathcal{D}^{M}_{k-1}\cup\mathcal{D}^{Z}_{k}$
3:  $\hat{X}^{M}_{0}\leftarrow X^{M}_{k}$
4:  $\hat{\theta}_{0}\leftarrow\theta_{k}$
@@
8:  $\mathcal{D}^{M}_{k}\leftarrow\mathcal{\hat{D}}_{n_{\alpha}}$
9:  $\theta_{k}\leftarrow\hat{\theta}_{n_{\alpha}}$
@@ III-A Trajectory Tracking
The primary aim of this paper is to provide precise tracking of reference trajectories in the presence of modeling uncertainties. To outline the trajectory tracking problem, we first define the reference trajectories. The path is expressed as a two-dimensional spline curve $\psi(s^{\mathrm{ref}}(t))$, defined by the coordinate functions $(x(s^{\mathrm{ref}}(t)),y(s^{\mathrm{ref}}(t))$, where $s^{\mathrm{ref}}$ is a time domain signal defined as $s^{\mathrm{ref}}:\mathbb{R}\rightarrow[0,L]$. The arc length of the full path is denoted as $L$, hence $s^{\mathrm{ref}}(t)$ describes the desired vehicle position along the path at time $t$. Both $x(s^{\mathrm{ref}}(t))$, and $y(s^{\mathrm{ref}}(t))$ are monotonic in $s^{\mathrm{ref}}(t)$, furthermore $(x(0),y(0))$ and $(x(L),y(L))$ assign the endpoints of the curve. The speed reference $v^{\mathrm{ref}}(s^{\mathrm{ref}}(t))=v^{\mathrm{ref}}(t)$ along the trajectory is also given. These types of reference motion trajectories can be obtained by regular path planning algorithms.
@@ III-B Baseline Vehicle Model
$\dot{x}$ $=v_{\xi}\cos(\varphi)-v_{\eta}\sin(\varphi),$ (12a)
$\dot{y}$ $=v_{\xi}\sin(\varphi)+v_{\eta}\cos(\varphi),$ (12b)
@@
$\dot{v}_{\xi}$ $=\frac{1}{m}\left(F_{\xi}+F_{\xi}\cos(\delta)-F_{\mathrm{f,\eta}}\sin(\delta)+mv_{\eta}\omega\right),$ (12d)
$\dot{v}_{\eta}$ $=\frac{1}{m}\left(F_{\mathrm{r,\eta}}+F_{\xi}\sin(\delta)+F_{\mathrm{f,\eta}}\cos(\delta)-mv_{\xi}\omega\right),$ (12e)
$\dot{\omega}$ $=\frac{1}{I_{z}}\left(F_{\mathrm{f,\eta}}l_{\mathrm{f}}\cos(\delta)+F_{\mathrm{f,\xi}}l_{\mathrm{f}}\sin(\delta)-F_{\mathrm{r,\eta}}l_{\mathrm{r}}\right),$ (12f)
where $(x,y)$ is the position and $\varphi$ is the orientation of the vehicle in the global coordinate frame. The states $v_{\xi}$ and $v_{\eta}$ denote the longitudinal and lateral velocity of the vehicle in a body fixed frame and $\omega$ is the yaw rate. The parameters of the model are the distance of the front and rear axis from the center of mass, denoted as $l_{\mathrm{f}}$ and $l_{\mathrm{r}}$, the mass of the vehicle $m$ and the inertia along the vertical axis $I_{z}$.
@@
The longitudinal tire force $F_{\xi}$ is determined by a drivetrain model, which assumes a first-order connection between the motor input and the velocity of the vehicle. This modeling technique has been successfully utilized previously in [24] and [4] for electric vehicles, hence, we adopted the following variant:
$F_{\xi}=C_{\mathrm{m1}}d-C_{\mathrm{m2}}v_{\xi}-C_{\mathrm{m3}},$ (13)
@@
$F_{\mathrm{r,\eta}}$ $=C_{\mathrm{r}}\arctan\left((-v_{\eta}+l_{\mathrm{r}}\omega)/v_{\xi}\right),$ (14a)
$F_{\mathrm{f,\eta}}$ $=C_{\mathrm{f}}\arctan\left(\delta-(v_{\eta}+l_{\mathrm{f}}\omega)/{v_{\xi}}\right),$ (14b)
where $C_{\mathrm{f}}$ and $C_{\mathrm{r}}$ are the cornering stiffness of the front and rear tire respectively. Finally, the control inputs of the vehicle are the steering angle $\delta$ and the motor input $d$, which a controller can directly actuate.
@@
$\dot{s}$ $=(v_{\xi}\cos(\theta_{\mathrm{e}})-v_{\eta}\sin(\theta_{\mathrm{e}}))/(1-c(s)e_{s}),$ (15a)
$\dot{e}_{s}$ $=v_{\xi}\sin(\theta_{\mathrm{e}})+v_{\eta}\cos(\theta_{\mathrm{e}}),$ (15b)
$\dot{\theta}_{\mathrm{e}}$ $=\omega-c(s)\dot{s},$ (15c)
where the newly introduced states are the position $s$ along the path, the lateral deviation $e_{s}$ and the heading error $\theta_{\mathrm{e}}$, while the lateral ($v_{\eta}$), longitudinal ($v_{\xi}$) and angular velocities ($\omega$) are the same as in (12). Furthermore, $c(s)$ describes the curvature of the reference path at $s$.
@@ IV-B Decoupling
$\dot{s}$ $=(v_{\xi}\cos(\theta_{\mathrm{e}})-v_{\eta}\sin(\theta_{\mathrm{e}}))/({1-c(s)e_{s}}),$ (16a)
@@
$\begin{split}\dot{v}_{\xi}&=\frac{1}{m}\big{(}(1+\cos(\delta))(C_{\mathrm{m1}}d-C_{\mathrm{m2}}v_{\xi}-C_{\mathrm{m3}})\\ &\qquad-F_{\mathrm{f,\eta}}\sin(\delta)+mv_{\eta}\omega\big{)},\end{split}$ (16b)
@@
where the state vector $x_{\mathrm{lo}}=[s\;v_{\xi}]^{\top}$ is the position along the path and the longitudinal velocity, while $d$ is the actuated control input. Furthermore, the heading error $\theta_{\mathrm{e}}$, lateral deviation $e_{s}$, lateral velocity $v_{\eta}$, steering angle $\delta$ and $F_{\mathrm{f,\eta}}$ are considered as external varying parameters depending on the lateral subsystem. The reference input of the system is $x_{\mathrm{lo}}^{\mathrm{ref}}=[s^{\mathrm{ref}}\;v^{\mathrm{ref}}]^{\top}$ defined by the trajectory. To simplify the control design and analysis, we separate the position and velocity states by introducing a virtual velocity generator that provides a position-adjusted virtual reference velocity as $v^{\mathrm{r}}=v^{\mathrm{ref}}-k_{v}(s-s^{\mathrm{ref}})$. This results in the modified control objective $v_{\xi}\rightarrow v^{\mathrm{r}}$, while the model dimensions are reduced, as only the dynamics of $v_{\xi}$ are needed to be considered. Therefore, the longitudinal behavior can be expressed as:
$\underbrace{\dot{v}_{\xi}}_{\dot{\raisebox{2.0pt}{$\chi$}}_{\mathrm{lo}}}=$ $\underbrace{\frac{-C_{\mathrm{m2}}(1+\cos(\delta))}{m}}_{A_{\mathrm{lo}}(\delta)}\underbrace{v_{\xi}}_{\raisebox{2.0pt}{$\chi$}_{\mathrm{lo}}}+\underbrace{\frac{C_{\mathrm{m1}}(1+\cos(\delta))}{m}}_{B_{\mathrm{lo}}(\delta)}\underbrace{d}_{u_{\mathrm{lo}}}$
$+\underbrace{\frac{-C_{\mathrm{m3}}(1+\cos(\delta))}{m}}_{w_{0}}+\underbrace{\frac{-F_{\mathrm{f,\eta}}\sin(\delta)}{m}+v_{\eta}\omega}_{w_{1}}$ (17)
@@
where $A_{\mathrm{lo}}(\delta)$ and $B_{\mathrm{lo}}(\delta)$ can be considered as parameter-varying state-transition and input matrices with steering angle $\delta$ as a scheduling variable, resulting in a linear parameter-varying (LPV) embedding [25]. Here, $w_{0}$ is a nonlinearity introduced by the dry friction of the drivetrain and $w_{1}$ lumps together the effects of the lateral subsystem.
From (15), the lateral behavior of the vehicle is described as
$\dot{e}_{s}$ $=v_{\xi}\sin(\theta_{\mathrm{e}})+v_{\eta}\cos(\theta_{\mathrm{e}}),$ (18a)
$\dot{\theta}_{\mathrm{e}}$ $=\omega-c(s)\dot{s},$ (18b)
$\dot{v}_{\eta}$ $=\frac{1}{m}\left(F_{\mathrm{r,\eta}}+F_{\xi}\sin(\delta)+F_{\mathrm{f,\eta}}\cos(\delta)-mv_{\xi}\omega\right),$ (18c)
$\dot{\omega}$ $=\frac{1}{I_{z}}\left(F_{\mathrm{f,\eta}}l_{\mathrm{f}}\cos(\delta)+F_{\mathrm{f,\xi}}l_{\mathrm{f}}\sin(\delta)-F_{\mathrm{r,\eta}}l_{\mathrm{r}}\right),$ (18d)
@@
where the state vector $x_{\mathrm{la}}=[e_{s}\;\theta_{\mathrm{e}}\;v_{\eta}\;\omega]^{\top}$ consists of the lateral error $e_{s}$, heading error $\theta_{\mathrm{e}}$, lateral velocity $v_{\eta}$ and yaw rate $\omega$, while $v_{\xi}$ is a scheduling variable. The input is steering angle $\delta$. To simplify the model, we first substitute the lateral tire models into (18), use small angle approximations ($\sin(\alpha)\approx\alpha$, $\cos(\alpha)\approx 1$), neglect the longitudinal tire force ($F_{\xi}\approx 0$) and approximate the velocity along the path as $\dot{s}\approx v_{\xi}$ which leads to the model:
@@
$\begin{split}\begin{bmatrix}\dot{e}_{s}\\ \dot{v}_{\eta}\\ \dot{\theta}_{\mathrm{e}}\\ \dot{\omega}\end{bmatrix}&=\begin{bmatrix}0&1&v_{\xi}&0\\ 0&A_{vv}&0&A_{v\omega}\\ 0&0&0&1\\ 0&A_{\omega v}&0&A_{\omega\omega}\end{bmatrix}\!\!\!\begin{bmatrix}{e}_{s}\\ v_{\eta}\\ \theta_{\mathrm{e}}\\ \omega\end{bmatrix}+\begin{bmatrix}0\\ \frac{C_{\mathrm{f}}}{m}\\ 0\\ \frac{l_{\mathrm{f}}C_{\mathrm{f}}}{I_{z}}\end{bmatrix}\delta+\begin{bmatrix}0\\ 0\\ v_{\xi}\\ 0\end{bmatrix}c(s),\end{split}$ (19)
@@
where $A_{vv}=-\frac{C_{\mathrm{f}}+C_{\mathrm{r}}}{mv_{\xi}}$, $A_{v\omega}=-v_{\xi}-\frac{l_{\mathrm{f}}C_{\mathrm{f}}-l_{\mathrm{r}}C_{\mathrm{r}}}{mv_{\xi}}$, $A_{\omega v}=\frac{l_{\mathrm{r}}C_{\mathrm{r}}-l_{\mathrm{f}}C_{\mathrm{f}}}{I_{z}v_{\xi}}$, $A_{\omega\omega}=-\frac{l_{\mathrm{f}}^{2}C_{\mathrm{f}}+l_{\mathrm{r}}^{2}C_{\mathrm{r}}}{I_{z}v_{\xi}}$. Note, that in (19) the path curvature is regarded as an external disturbance. Further simplification can be achieved by expressing the lateral model only in terms of the error variables and their derivatives, i.e., by introducing $\ddot{e}_{s}=\dot{v}_{\eta}+v_{\xi}\dot{\theta}_{\mathrm{e}}$. As discussed in [26], regulating both the heading and the lateral error to the origin yields poor tracking performance, as the two quantities cannot be simultaneously zero along the path if we assume perfect tracking. Therefore, we separate the lateral error and the heading dynamics. We use the lateral error dynamics for feedback control design and only regulate the heading error with a feedforward, similar to the Stanley controller [27]. Furthermore, we also introduce the integral of the lateral error, i.e., $q(t)=\int_{0}^{t}e_{s}(\tau)\mathrm{d}\tau$ as a new state to assure asymptotic convergence of $e_{s}$ without offset. The final control-oriented lateral model can be expressed as

### v2

@@ II-C Online Learning and Update for SGPs
The VFE method is constructed for offline training, i.e., it requires the complete training set to be available. However, to provide adaptivity to changing environmental conditions, online training is necessary, which updates the model during operation using the latest measurement data. For training GPs online, various algorithms such as the recursive least squares GP (RLS-GP) [11] or the dynamic sparse GP (DS-GP) [10] are available. This article proposes a recursive gradient-based (RGB) optimization scheme for SGPs utilizing the VFE method outlined in Sec. II-B. Compared to the previously introduced techniques that either update the approximate posterior distribution (RLS-GP) or the inducing points (DS-GP), this approach jointly updates both the hyperparameters and the inducing points of the GP based on the incoming new information, which, to the authors’ knowledge, has not been applied before.
First, we assume that the SGP has been trained offline on the initial dataset $\mathcal{D}^{N}$, yielding the optimal values for $\theta$ hyperparameters and $X^{M}$ inducing points that correspond to $Y^{M}$ virtual outputs. As outlined before, the core idea of sparse GP estimation is that the virtual $\mathcal{D}^{M}$ dataset is obtained such that its distribution approximates the original $\mathcal{D}^{N}$ set with considerably fewer data points. Therefore, in future predictions, we can rely only on the $\mathcal{D}^{M}$ set, as it represents the information content of $\mathcal{D}^{N}$.
Let the one incoming batch of input-output data at time step $k$ be denoted as $\mathcal{D}^{Z}_{k}=\{X^{Z}_{k},Y^{Z}_{k}\}$, where $X^{Z}_{k}=[{{\color[rgb]{0,0,0}\mathpzc{x}}}_{1}^{\top}\cdots{{\color[rgb]{0,0,0}\mathpzc{x}}}_{Z}^{\top}]\in\mathbb{R}^{Z\times n_{\mathrm{x}}}$, $Y^{Z}_{k}=[{{\color[rgb]{0,0,0}\mathpzc{y}}}_{1}\cdots{{\color[rgb]{0,0,0}\mathpzc{y}}}_{Z}]\in\mathbb{R}^{Z}$ and $Z$ is the batch size. Then, after receiving one batch of new training data at time $k$, we can combine the previously obtained pseudo dataset and the new batch to define a new training dataset as $\mathcal{\hat{D}}_{k}=\mathcal{D}^{M}_{k-1}\cup\mathcal{D}^{Z}_{k}$, where $\mathcal{D}_{0}^{M}=\mathcal{D}^{M}$. Note that with this formulation, $\mathcal{\hat{D}}_{k}$ both contains information from the pre-trained GP and innovation from the update batch. By substituting this combined dataset into the VFE cost (8), we can perform $n_{\alpha}$ steps of gradient descent to collectively update the hyperparameters of the GP and the inducing points, i.e.
@@
where $\alpha$ is the learning rate and $i\in\mathbb{I}_{1}^{n_{\alpha}}$. Note that the evaluation of (11) has the computational complexity of $\mathcal{O}((M+Z)M^{2})$, therefore, gradient descent can be executed multiple times online in one update iteration. Furthermore, using the augmented training dataset $\mathcal{\hat{D}}_{k}$ is beneficial as it incorporates both the previously learned knowledge from $D^{M}_{k-1}$ and the new information contained in $\mathcal{D}_{k}^{Z}$. The RBG algorithm is outlined in Alg. 1 and has three important parameters: size of the pseudo dataset $M$, update batch size $Z$ and $n_{\alpha}$ number of retrain iterations and $\alpha$ learning rate. The learning rate can be used to determine how much the GPs should favor new incoming data, i.e. how fast the GPs adapt to changing conditions. However, note that too fast adaptation can lead to forgetting, similarly to the RSL-GP.
@@
[…]
@@
8: Update $\hat{\mathcal{D}}^{M}_{k}$ with $\hat{X}_{n_{\alpha}}^{M}$
[…]
@@ III-A Trajectory Tracking
The primary aim of this paper is to provide precise tracking of reference trajectories in the presence of modeling uncertainties. To outline the trajectory tracking problem, we first define the reference trajectories. The path is expressed as a two-dimensional spline curve $\psi(s^{\mathrm{ref}}(t))$, defined by the coordinate functions $(x(s^{\mathrm{ref}}(t)),y(s^{\mathrm{ref}}(t))$, where $s^{\mathrm{ref}}$ is a time domain signal defined as $s^{\mathrm{ref}}:\mathbb{R}\rightarrow[0,L]$. The arc length of the full path is denoted as $L$, hence $s^{\mathrm{ref}}(t)$ describes the desired vehicle position along the path at time $t$. Both $x(s^{\mathrm{ref}}(t))$, and $y(s^{\mathrm{ref}}(t))$ are continuous in $s^{\mathrm{ref}}(t)$, furthermore $(x(0),y(0))$ and $(x(L),y(L))$ assign the endpoints of the curve. The speed reference $v^{\mathrm{ref}}(s^{\mathrm{ref}}(t))=v^{\mathrm{ref}}(t)$ along the trajectory is also given. These types of reference motion trajectories can be obtained by regular path planning algorithms.
@@ III-B Baseline Vehicle Model
${{\color[rgb]{0,0,0}\dot{p}_{\mathrm{x}}}}$ $=v_{\mathrm{{{\color[rgb]{0,0,0}x}}}}\cos(\varphi)-v_{\mathrm{{{\color[rgb]{0,0,0}y}}}}\sin(\varphi),$ (12a)
${{\color[rgb]{0,0,0}\dot{p}_{\mathrm{y}}}}$ $=v_{\mathrm{{{\color[rgb]{0,0,0}x}}}}\sin(\varphi)+v_{\mathrm{{{\color[rgb]{0,0,0}y}}}}\cos(\varphi),$ (12b)
@@
$\dot{v}_{\mathrm{{{\color[rgb]{0,0,0}x}}}}$ $=\frac{1}{m}\left(F_{\mathrm{{{\color[rgb]{0,0,0}x}}}}+F_{\mathrm{x}}\cos(\delta)-F_{\mathrm{f,{{\color[rgb]{0,0,0}y}}}}\sin(\delta)+mv_{\mathrm{{{\color[rgb]{0,0,0}y}}}}\omega\right),$ (12d)
$\dot{v}_{\mathrm{{{\color[rgb]{0,0,0}y}}}}$ $=\frac{1}{m}\left(F_{\mathrm{r,{{\color[rgb]{0,0,0}y}}}}+F_{\mathrm{{{\color[rgb]{0,0,0}x}}}}\sin(\delta)+F_{\mathrm{f,{{\color[rgb]{0,0,0}y}}}}\cos(\delta)-mv_{\mathrm{{{\color[rgb]{0,0,0}x}}}}\omega\right),$ (12e)
$\dot{\omega}$ $=\frac{1}{I_{\mathrm{z}}}\left(F_{\mathrm{f,{{\color[rgb]{0,0,0}y}}}}l_{\mathrm{f}}\cos(\delta)+F_{\mathrm{{{\color[rgb]{0,0,0}x}}}}l_{\mathrm{f}}\sin(\delta)-F_{\mathrm{r,{{\color[rgb]{0,0,0}y}}}}l_{\mathrm{r}}\right),$ (12f)
[…]
@@
The longitudinal tire force $F_{\mathrm{x}}$ is computed using a drivetrain model that assumes a first-order relationship between the motor input and the longitudinal velocity. This modeling approach has been effectively applied to electric vehicles in prior work [25, 3], therefore, we adopt a similar variant in this study:
$F_{\mathrm{{{\color[rgb]{0,0,0}x}}}}=C_{\mathrm{m1}}d-C_{\mathrm{m2}}v_{\mathrm{{{\color[rgb]{0,0,0}x}}}}-C_{\mathrm{m3}},$ (13)
@@
$F_{\mathrm{r,{{\color[rgb]{0,0,0}y}}}}$ $=C_{\mathrm{r}}\arctan\left((-v_{\mathrm{{{\color[rgb]{0,0,0}y}}}}+l_{\mathrm{r}}\omega)/v_{\mathrm{{{\color[rgb]{0,0,0}x}}}}\right),$ (14a)
$F_{\mathrm{f,{{\color[rgb]{0,0,0}y}}}}$ $=C_{\mathrm{f}}\arctan\left(\delta-(v_{\mathrm{{{\color[rgb]{0,0,0}y}}}}+l_{\mathrm{f}}\omega)/{v_{\mathrm{{{\color[rgb]{0,0,0}x}}}}}\right),$ (14b)
where $C_{\mathrm{f}}$ and $C_{\mathrm{r}}$ are the cornering stiffness of the front and rear tire, respectively. Finally, the control inputs of the vehicle are the steering angle $\delta$ and the motor input $d$, which a controller can directly actuate.
@@
$\dot{s}$ $=(v_{\mathrm{{{\color[rgb]{0,0,0}x}}}}\cos(\theta_{\mathrm{e}})-v_{\mathrm{{{\color[rgb]{0,0,0}y}}}}\sin(\theta_{\mathrm{e}}))/(1-{{\color[rgb]{0,0,0}\kappa}}(s)e_{s}),$ (15a)
$\dot{e}_{s}$ $=v_{\mathrm{{{\color[rgb]{0,0,0}x}}}}\sin(\theta_{\mathrm{e}})+v_{\mathrm{{{\color[rgb]{0,0,0}y}}}}\cos(\theta_{\mathrm{e}}),$ (15b)
$\dot{\theta}_{\mathrm{e}}$ $=\omega-{{\color[rgb]{0,0,0}\kappa}}(s)\dot{s},$ (15c)
where the newly introduced states are the position $s$ along the path, the lateral deviation $e_{s}$ and the heading error $\theta_{\mathrm{e}}$, while the lateral ($v_{\mathrm{{{\color[rgb]{0,0,0}y}}}}$), longitudinal ($v_{\mathrm{{{\color[rgb]{0,0,0}x}}}}$) and angular velocities ($\omega$) are the same as in (12). Furthermore, ${{\color[rgb]{0,0,0}\kappa}}(s)$ describes the curvature of the reference path at $s$.
@@ IV-B Decoupling
$\dot{s}$ $=(v_{\mathrm{{{\color[rgb]{0,0,0}x}}}}\cos(\theta_{\mathrm{e}})-v_{\mathrm{{{\color[rgb]{0,0,0}y}}}}\sin(\theta_{\mathrm{e}}))/({1-{{\color[rgb]{0,0,0}\kappa}}(s)e_{s}}),$ (16a)
@@
$\begin{split}\dot{v}_{\mathrm{{{\color[rgb]{0,0,0}x}}}}&=\frac{1}{m}\big((1+\cos(\delta))(C_{\mathrm{m1}}d-C_{\mathrm{m2}}v_{\mathrm{{{\color[rgb]{0,0,0}x}}}}-C_{\mathrm{m3}})\\ &\qquad-F_{\mathrm{f,{{{\color[rgb]{0,0,0}y}}}}}\sin(\delta)+mv_{\mathrm{{{\color[rgb]{0,0,0}y}}}}\omega\big),\end{split}$ (16b)
@@
where the state vector $x_{\mathrm{lo}}=[s\;v_{\mathrm{{{\color[rgb]{0,0,0}x}}}}]^{\top}$ is the position along the path and the longitudinal velocity, while $d$ is the actuated control input. Furthermore, the heading error $\theta_{\mathrm{e}}$, lateral deviation $e_{s}$, lateral velocity $v_{\mathrm{{{\color[rgb]{0,0,0}y}}}}$, steering angle $\delta$ and $F_{\mathrm{f,{{\color[rgb]{0,0,0}y}}}}$ are considered as external varying parameters depending on the lateral subsystem. The reference input of the system is $x_{\mathrm{lo}}^{\mathrm{ref}}=[s^{\mathrm{ref}}\;v^{\mathrm{ref}}]^{\top}$ defined by the trajectory. To simplify the control design and analysis, we separate the position and velocity states by introducing a virtual velocity generator that provides a position-adjusted virtual reference velocity as $v^{\mathrm{r}}=v^{\mathrm{ref}}-k_{v}(s-s^{\mathrm{ref}})$. This results in the modified control objective $v_{\mathrm{{{\color[rgb]{0,0,0}x}}}}\rightarrow v^{\mathrm{r}}$, while the model dimensions are reduced, as only the dynamics of $v_{\mathrm{{{\color[rgb]{0,0,0}x}}}}$ need to be considered. Therefore, the longitudinal behavior can be expressed as:
$\underbrace{\dot{v}_{\mathrm{{{\color[rgb]{0,0,0}x}}}}}_{\dot{\raisebox{2.0pt}{$\chi$}}_{\mathrm{lo}}}=$ $\underbrace{\frac{-C_{\mathrm{m2}}(1+\cos(\delta))}{m}}_{A_{\mathrm{lo}}(\delta)}\underbrace{v_{\mathrm{{{\color[rgb]{0,0,0}x}}}}}_{\raisebox{2.0pt}{$\chi$}_{\mathrm{lo}}}+\underbrace{\frac{C_{\mathrm{m1}}(1+\cos(\delta))}{m}}_{B_{\mathrm{lo}}(\delta)}\underbrace{d}_{u_{\mathrm{lo}}}$
$+\underbrace{\frac{-C_{\mathrm{m3}}(1+\cos(\delta))}{m}}_{w_{0}}+\underbrace{\frac{-F_{\mathrm{f,\eta}}\sin(\delta)}{m}+v_{\mathrm{{{\color[rgb]{0,0,0}y}}}}\omega}_{w_{1}}$ (17)
@@
where $A_{\mathrm{lo}}(\delta)$ and $B_{\mathrm{lo}}(\delta)$ can be considered as parameter-varying state-transition and input matrices with steering angle $\delta$ as a scheduling variable, resulting in a linear parameter-varying (LPV) embedding [26]. In (IV-B), $w_{0}$ is a nonlinearity introduced by the dry friction of the drivetrain and $w_{1}$ lumps together the effects of the lateral subsystem.
From (15), the lateral behavior of the car is described as
$\dot{e}_{s}$ $=v_{\mathrm{{{\color[rgb]{0,0,0}x}}}}\sin(\theta_{\mathrm{e}})+v_{\mathrm{{{\color[rgb]{0,0,0}y}}}}\cos(\theta_{\mathrm{e}}),$ (18a)
$\dot{\theta}_{\mathrm{e}}$ $=\omega-{{\color[rgb]{0,0,0}\kappa}}(s)\dot{s},$ (18b)
$\dot{v}_{\mathrm{{{\color[rgb]{0,0,0}y}}}}$ $=\frac{1}{m}\left(F_{\mathrm{r,{{\color[rgb]{0,0,0}y}}}}+F_{\mathrm{{{\color[rgb]{0,0,0}x}}}}\sin(\delta)+F_{\mathrm{f,{{\color[rgb]{0,0,0}y}}}}\cos(\delta)-mv_{\mathrm{{{\color[rgb]{0,0,0}x}}}}\omega\right),$ (18c)
$\dot{\omega}$ $=\frac{1}{I_{\mathrm{z}}}\left(F_{\mathrm{f,y}}l_{\mathrm{f}}\cos(\delta)+F_{\mathrm{{{\color[rgb]{0,0,0}x}}}}l_{\mathrm{f}}\sin(\delta)-F_{\mathrm{r,{{\color[rgb]{0,0,0}y}}}}l_{\mathrm{r}}\right),$ (18d)
@@
where the state vector $x_{\mathrm{la}}=[e_{s}\;\theta_{\mathrm{e}}\;v_{\mathrm{{{\color[rgb]{0,0,0}y}}}}\;\omega]^{\top}$ consists of the lateral error $e_{s}$, heading error $\theta_{\mathrm{e}}$, lateral velocity $v_{\mathrm{{{\color[rgb]{0,0,0}y}}}}$ and yaw rate $\omega$, while $v_{\mathrm{{{\color[rgb]{0,0,0}x}}}}$ is a scheduling variable. The input is steering angle $\delta$. To simplify the model, we first substitute the lateral tire models into (18), use small angle approximations ($\sin(\alpha)\approx\alpha$, $\cos(\alpha)\approx 1$), neglect the longitudinal tire force ($F_{\mathrm{{{\color[rgb]{0,0,0}x}}}}\approx 0$) and approximate the velocity along the path as $\dot{s}\approx v_{\mathrm{{{\color[rgb]{0,0,0}x}}}}$ which leads to the model [27]:
@@
$\begin{split}\begin{bmatrix}\dot{e}_{s}\\ \dot{v}_{\mathrm{{{\color[rgb]{0,0,0}y}}}}\\ \dot{\theta}_{\mathrm{e}}\\ \dot{\omega}\end{bmatrix}&=\begin{bmatrix}0&1&v_{\xi}&0\\ 0&A_{vv}&0&A_{v\omega}\\ 0&0&0&1\\ 0&A_{\omega v}&0&A_{\omega\omega}\end{bmatrix}\!\!\!\begin{bmatrix}{e}_{s}\\ v_{\mathrm{{{\color[rgb]{0,0,0}y}}}}\\ \theta_{\mathrm{e}}\\ \omega\end{bmatrix}+\begin{bmatrix}0\\ \frac{C_{\mathrm{f}}}{m}\\ 0\\ \frac{l_{\mathrm{f}}C_{\mathrm{f}}}{I_{z}}\end{bmatrix}\delta+\begin{bmatrix}0\\ 0\\ v_{\mathrm{{{\color[rgb]{0,0,0}x}}}}\\ 0\end{bmatrix}{{\color[rgb]{0,0,0}\kappa}}(s),\end{split}$ (19)
@@
where $A_{vv}=-\frac{C_{\mathrm{f}}+C_{\mathrm{r}}}{mv_{\mathrm{{{\color[rgb]{0,0,0}x}}}}}$, $A_{v\omega}=-v_{\mathrm{{{\color[rgb]{0,0,0}x}}}}-\frac{l_{\mathrm{f}}C_{\mathrm{f}}-l_{\mathrm{r}}C_{\mathrm{r}}}{mv_{\mathrm{{{\color[rgb]{0,0,0}x}}}}}$, $A_{\omega v}=\frac{l_{\mathrm{r}}C_{\mathrm{r}}-l_{\mathrm{f}}C_{\mathrm{f}}}{I_{z}v_{\mathrm{{{\color[rgb]{0,0,0}x}}}}}$, $A_{\omega\omega}=-\frac{l_{\mathrm{f}}^{2}C_{\mathrm{f}}+l_{\mathrm{r}}^{2}C_{\mathrm{r}}}{I_{\mathrm{z}}v_{\mathrm{{{\color[rgb]{0,0,0}x}}}}}$. Note that in (19) the path curvature is regarded as an external disturbance. Further simplification can be achieved by expressing the lateral model only in terms of the error variables and their derivatives, i.e., by introducing $\ddot{e}_{s}=\dot{v}_{\mathrm{{{\color[rgb]{0,0,0}y}}}}+v_{\mathrm{{{\color[rgb]{0,0,0}x}}}}\dot{\theta}_{\mathrm{e}}$. As discussed in [28], regulating both the heading and the lateral error to the origin yields poor tracking performance, as the two quantities cannot be simultaneously zero along the path if we assume perfect tracking. Therefore, we separate the lateral error and the heading dynamics. We use the lateral error dynamics for feedback control design and only regulate the heading error with a feedforward, similar to the Stanley controller [29]. Furthermore, we also introduce the integral of the lateral error, i.e., $q(t)=\int_{0}^{t}e_{s}(\tau)\mathrm{d}\tau$ as a new state to assure asymptotic convergence of $e_{s}$ without offset. The final control-oriented lateral model can be expressed as